Fermi National Laboratory


Accelerator Update for March 7, 2005 - March 9, 2005

Monday March 7
The day shift began with Operations monitoring store 4024 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiments taking beam.

The Neutron Therapy Facility (NTF) requested beam at 8:43 AM for patient treatment.

At 9:37 AM, a Booster kicker (MKS05) tripped off. An expert had the kicker back on about ten minutes later.

At 11:25 AM, the Recycler lost its stash while conducting studies.

At 11:46 AM, a Booster RF station (BRF04) tripped off for the first of three times in twenty minutes. Operators bypassed the station after the last trip and called an expert.

A Booster expert had BRF04 back on by 12:10 PM, but then it tripped off again ten minutes later.

At 12:57 PM, the Main Injector (MI) power supply interlock tripped off. Operators contacted experts who investigated, but couldn't find anything wrong.

NTF finished using the beam at 1 PM.

Operations reported at 1:13 PM that they had MI back on line and that Pbar had resumed stacking.

Operations reported at 1:23 PM that the MI and NuMI experts were beginning to commission the new stacking/NuMI Time Line Generator (TLG) module.

At 1:32 PM, MI lost its beam permit due to a MECAR trip. Experts investigated the failure and had MI back up by 1:58 PM.

At 2:48 PM, a MI RF station (MIRF11) tripped off for the first of three times in an hour. It reset okay each time.

At 3:25 PM, two Linac RF stations (KRF4 & 6) tripped off. They both reset okay.

At 3:27 PM, KRF6 tripped off again. Operators turned off Linac to investigate.

At 4:42 PM, MIRF11 tripped off. It reset okay.

At 5:16 PM, Linac experts reported that they were replacing some bad capacitors on KRF6. They had this done and beam to Booster by 5:35 PM.

Operations reestablished beam to MiniBooNE at 5:40 PM.

Operations began stashing antiprotons into the Recycler at 5:46 PM.

Operations established beam to Meson MTest experiment at 6:54 PM

At 11:36 PM, a Booster interlock monitor had its power supply fail, which caused the Booster RF to trip off.

Tuesday March 8
The midnight shift began with Operations monitoring store 4024 and the stash, and with no downstream beam due to a Booster interlock power supply failure.

Operator replaced the Booster interlock crate power supply and resumed sending beam to MiniBooNE, Pbar, and MCenter at 1:20 AM.

MTest experimenters said they were done with the beam for the night at 1:30 AM.

Pbar experts ended their study and resumed stacking at 3:31 AM.

At 9:24 AM, the H- Source had an intensity drop. Operators contacted an expert who adjusted the source at 9:38 AM.

At 10:38 AM, BRF4 tripped off for the first of three times in two minutes. Operators contacted an expert.

Operations terminated store 4024 at 11:07 AM.

Cryo system technicians began maintenance on the Tevatron (TeV) sector C3 wet engine at 11:21 AM. This work took about 35 minutes.

A Pbar expert began a reverse proton study at 11:22 AM. A TeV expert loaded some beam corrections into the TeV at the same time.

Operators made the D-Zero (D0) collision hall ready for access at 11:25 AM. They were out of access an hour later.

Pbar studies ended at 1:27 PM.

Operations put the TeV into shot setup at 2:48 PM. Shot setup halts all NuMI and Switchyard (SY) beam.

A Preaccelerator expert began adjusting the H- Source at 5:02 PM. He worked on this for about ninety minutes.

Operations established store 4027 at 5:03 PM with an initial luminosity of 93.36E30.

At 5:09 PM, the TeV suffered a multi-house quench due to an antiproton kicker prefire.

Operations tried to send beam to MTest at 5:55 PM, but closed vacuum valves in the P3 transfer line to SY hindered the attempt. Operators investigated and then sent beam.

Operators prepared the Collider Detector at Fermilab (CDF) for access at 5:57 PM.

At 6:35 PM, Operations reported that the P3 transfer line vacuum valves showed closed again, but since beam was on at the time, operators could see the beam.

At 6:34 PM, Operations reported the output from the H- Source dropped again.

At 6:52 PM, Linac's drift tube 3 tripped off for the first of six times in about twenty minutes. Operations reported that this was due to turning on NTF and to adjusting the H- Source.

At 8:53 PM, Operations reported problems with stacking due to an Accumulator RF station (ARF1).

Operations established beam to MCenter at 9:54 PM.

At 10:30 PM, a Debuncher kicker (D:EKIK) tripped off and wouldn't reset. Operators called in an expert.

NTF said that they were done with the beam at 11:35 PM.

Wednesday March 9
The midnight shift began with operations monitoring the stack and the stash, with CDF in access, with an expert investigating a Debuncher kicker, and with MiniBooNE and MCenter taking beam.

MTest experimenters said that they were through with the beam for the night at 12:12 AM.

The Debuncher kicker work finished and Pbar resumed stacking at 12:54 AM.

At 1:24 AM, Operations was trying to set up the TeV for studies when a problem with a correction magnet (T:HA48) developed. Operators changed out the power supply, but that didn't help. Operators discovered that the magnet was bad. They bypassed the magnet and a TeV expert tuned around the magnet.

TeV studies ended at 6:20 AM.

Operations put the TeV into shot setup at 7:08 AM.

The Plan for Monday
The plan for today is to establish a new store and then resume stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/9/2005   email Fermilab